What would be the new volume if the pressure on 600 mL is increased from 90
kPa to 150 kPa?
1
Expert's answer
2020-05-19T08:45:03-0400

According to the equations of Mendeleev-Clapeyron for initial and final states:

"p(1)•V(1)=\nu•R•T"

"p(2)•V(2)=\nu•R•T"

We can divide the first equation into the second equation and then we will get this:

"p(1)\u2022V(1)=p(2)\u2022V(2)"

So, "V(2)=(p(1)\u2022V(1))\/p(2)" = (90(kPa)•600(ml))/150(kPa) = 360 ml

Answer: V(2) = 360 ml